OpenStudy (anonymous):

Sylvia and Leonard are swimming laps in the pool and running laps around the track. The equations below can be used to determine how much of each type of exercise Sylvia and Leonard can complete in one hour, where x represents the number of laps completed in the pool and y represents the number of laps completed around the track. Total number of pool laps and track laps made by Sylvia: x + y = 25 Total number of pool laps and track laps made by Leonard: 2x + 3y = 60 When Sylvia and Leonard each swim 15 laps in the pool, what is the total number of laps around the track they can run?

OpenStudy (anonymous):

substitute 15 for x in each equation and solve for y: x + y = 25 15 + y = 25 2x + 3y = 60 2(15) + 3y = 60

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

any time!